A rectangular box has total surface area
600 cm^2 and a square base with edge length x cm.Express the volume V of the box as a function of x. What is the volume of the box if the edge length is 5 cm?
1
Expert's answer
2020-09-28T08:49:20-0400

Let x be the length of the side of the square base a h the height of the parallelepiped

then volume

"V=x*x*h"

S =600  total surface area

"S = 2*x^2 + 4*x*h"

"h = \frac{S-2*x^2}{4*x} = \frac{600-2*x^2}{4*x}"

"V = \frac{600-2*x^2}{4*x} *x^2 = 150*x -\frac{x^3}{2}"

for x = 5 cm V = 150*5 -125/2 = 687.5 cm3


Answer : "V = 150*x -\frac{x^3}{2}" ; for x = 5 cm V = 150*5 -125/2 = 687.5 cm3